You propose to build a space ship that allows the entire population of the United States (280 million) to leave the earth. Your plans allocate an average space of 1000=10X10X10 cubic feet to each person. Your space ship is going to have the shape of a sphere. The radius of your space ship is_____feet.

Answer by kapilsinghi(68)

Space required by one person = 1000 CF
Space required by 280,000,000 peopel = 280,000,000,000 CF
Let the radius of the spher is r
so volume of the sphere = Pi*r^3
volume of people should be equal to the volume of sphere
so Pi*r^3 = 280,000,000,000
r^3 = 280,000,000,000/3.14
r^3 = (280/3.14)* 1000,000,000
r^3 = 89*1000,000,000
r = (89 * 1000,000,000)^1/3
= (89)^1/3 * 1000
r = 4.4 * 1000 Feet

kapilsinghi123@gmail.com